# Effects of personality steering on cooperative behavior in Large Language Model agents

### v1

@@ 3.2 Comparison between baseline and personality-informed conditions
Figure 1: Average cooperation rates (left) and average cumulative payoffs (right) under the baseline condition (dashed) and the personality-informed condition (solid).
@@
For GPT-3.5-turbo, cooperation rates against non-cooperative strategies such as ALLD and RANDOM were relatively high even in the baseline condition, compared to GPT-4o and GPT-5. Under the personality-informed condition, cooperation rates against these strategies increased substantially, while average cumulative payoffs decreased. This result indicates that although explicit personality information enhances cooperation, it also increases vulnerability to exploitation by defecting or unpredictable opponents. In contrast, cooperation rates and payoffs against cooperative or reciprocal strategies, ALLC, GRIM TRIGGER, and TFT, remained high and largely unchanged across both conditions.
GPT-4o exhibited lower cooperation rates against ALLD and RANDOM strategies in the baseline condition, indicating more cautious behavior toward non-cooperative opponents. However, under the personality-informed condition, cooperation rates against these strategies increased, accompanied by a reduction in cumulative payoffs. Notably, the increase in cooperation was more pronounced for RANDOM than for ALLD, suggesting that GPT-4o distinguishes between purely defecting strategies and those that allow for potential reciprocal cooperation.
For GPT-5, baseline behavior against ALLD and RANDOM was broadly similar to that of GPT-4o, whereas cooperation rates against ALLC, GRIM TRIGGER, and TFT were slightly lower than those observed for the other models. An inspection of action histories reveals that GPT-5 frequently defected in the final round while maintaining cooperation in earlier rounds, a pattern consistent with end-game optimization in finitely repeated games. Explanations generated by the model often refer to balancing long-term trust-building with payoff maximization in the final round, suggesting a more strategic decision-making process. Notably, such payoff-maximizing and seemingly self-interested behavior is consistent with prior findings showing that models with explicit reasoning capabilities tend to exhibit more self-interested decision-making compared to models without explicit reasoning processes [29]. Under the personality-informed condition, cooperation rates against ALLC, GRIM TRIGGER, and TFT approached 100%, while cumulative payoffs decreased. Cooperation rates against ALLD remained largely unchanged, and cooperation against RANDOM increased only modestly, indicating that GPT-5 limited cooperation with exploitative opponents even when personality information promoted cooperativeness.
@@ 3.3 Effects of extreme personality manipulation
Finally, we report the results of Experiment 3, in which we artificially manipulated individual Big Five personality traits to extreme values. Using the average BFI scores obtained under the personality-informed condition in Experiment 1 as a reference, we independently replaced one personality dimension, openness, conscientiousness, extraversion, agreeableness, or neuroticism, with an extreme value of either 1 or 5, while keeping the remaining dimensions fixed. The RPD experiments were then conducted under the same settings as in Experiment 2. Figures 2, 3, and 4 present the average cooperation rates, average cumulative payoffs, and their differences for GPT-3.5-turbo, GPT-4o, and GPT-5, respectively.
@@
For GPT-3.5-turbo, manipulation of agreeableness had the strongest impact on cooperative behavior, followed by extraversion. When agreeableness was set to 1 (A-1), cooperation rates were consistently lower than those observed under other personality manipulations across all opponent strategies, resulting in a large negative difference in average cooperation. In contrast, when agreeableness was set to 5 (A-5), cooperation rates were comparable to those observed under other personality conditions, yielding relatively small differences.
Figure 2: Average cooperation rates and average cumulative payoffs for GPT-3.5-turbo under different personality score settings. The top row corresponds to setting the manipulated dimension to 1, the middle row to 5, and the bottom row shows the difference (5-1).
A similar but weaker pattern is observed for extraversion. Under the E-1 condition, cooperation rates against ALLD and RANDOM strategies were slightly lower than those under other personality manipulations, which contributes to observable differences in average cooperation. When extraversion is set to 5 (E-5), however, cooperation rates show little deviation from other conditions.
For GPT-4o, agreeableness again emerged as the dominant factor influencing cooperative behavior, whereas manipulation of other personality traits had minimal effects. Under the A-1 condition, cooperation rates droped to 0% across all opponent strategies, indicating consistently defecting behavior. Under the A-5 condition, average cooperation rates increased to levels comparable to or higher than those observed under other personality manipulations. However, cooperation rates against ALLD and RANDOM strategies were highest under A-5, resulting in increased exploitation and, consequently, the lowest average cumulative payoff among all personality conditions.
@@
Figure 3: Average cooperation rates (left) and average cumulative payoffs (right) for GPT-4o under extreme personality-score manipulations. The top, middle, and bottom rows are defined in the same way as in Fig. 2.
@@
For GPT-5, agreeableness similarly had the largest effect on cooperation, while manipulation of other personality traits produces little changed in behavior. Under the A-1 condition, cooperation rates fell to 0% across all opponent strategies, mirroring the behavior observed for GPT-4o. Under the A-5 condition, cooperation rates increased across strategies; however, the increase in cooperation against ALLD and RANDOM was smaller than that observed for GPT-4o. As a result, GPT-5 experienced less exploitation under high agreeableness than GPT-4o.
Figure 4: Average cooperation rates (left) and average cumulative payoffs (right) for GPT-5 under extreme personality-score manipulations. The top, middle, and bottom rows are defined in the same way as in Fig. 2.
@@ 4 Conclusion
In this study, we investigated how artificially controlled personality traits influence cooperative behavior in large language model agents. Using the Big Five framework, we first quantitatively measured basic personality profiles of multiple LLMs, and then examined their behavior in repeated Prisoner’s Dilemma games under conditions with and without explicit personality information. We further analyzed the effects of independently manipulating each personality dimension to extreme values.
@@
Taken together, our study indicates that personality steering alone is insufficient to fully control cooperative behavior in LLM agents. Rather, cooperative outcomes emerge from the interaction between personality cues and model-specific decision-making processes. This highlights the importance of considering both personality and strategic reasoning when analyzing and designing cooperative AI agents.

### v2

@@ 3.2 Comparison between baseline and personality-informed conditions
Figure 2: Average cooperation rates (left) and average cumulative payoffs (right) under the baseline condition (dashed) and the personality-informed condition (solid): (A) GPT-3.5-turbo, (B) GPT-4o, and (C) GPT-5.
@@
For GPT-3.5-turbo (Fig. 2A), cooperation rates against non-cooperative strategies such as ALLD and RANDOM were relatively high even in the baseline condition, compared to GPT-4o (Fig. 2B) and GPT-5 (Fig. 2C). Under the personality-informed condition, cooperation rates against these strategies increased substantially, while average cumulative payoffs decreased slightly. This result indicates that although explicit personality information enhances cooperation, it also increases vulnerability to exploitation by defecting or unpredictable opponents. In contrast, cooperation rates and payoffs against cooperative or reciprocal strategies, ALLC, GRIM TRIGGER, and TFT, remained high and largely unchanged across both conditions in GPT-3.5-turbo.
GPT-4o exhibited lower cooperation rates against ALLD and RANDOM strategies in the baseline condition, indicating more cautious behavior toward non-cooperative opponents (Fig. 2B). However, under the personality-informed condition, cooperation rates against these strategies increased, accompanied by a reduction in cumulative payoffs. Notably, the increase in cooperation was more pronounced for RANDOM than for ALLD, suggesting that GPT-4o distinguishes between purely defecting strategies and those that allow for potential reciprocal cooperation.
For GPT-5 (Fig. 2C), baseline behavior against ALLD was broadly similar to that of GPT-4o, whereas cooperation against RANDOM was substantially lower. In addition, cooperation rates against ALLC, GRIM TRIGGER, and TFT were lower than those observed for GPT-4o. By closely examining the game action histories, we found that GPT-5 frequently defected in the final round while maintaining cooperation in earlier rounds, a pattern consistent with end-game optimization in finitely repeated games. Notably, such payoff-maximizing and seemingly self-interested behavior is consistent with prior findings showing that models with explicit reasoning capabilities tend to exhibit more self-interested decision-making compared to models without explicit reasoning processes [Li2025]. Under the personality-informed condition, cooperation rates against ALLC, GRIM TRIGGER, and TFT approached 100%. Cooperation rates against ALLD remained largely unchanged, and cooperation against RANDOM increased only modestly, indicating that GPT-5 limited cooperation with exploitative opponents even when personality information promoted cooperativeness.
@@ 3.3 Effects of extreme personality manipulation
Finally, we report the results of Experiment 3, in which we artificially manipulated individual Big Five personality traits to extreme values. Using the average BFI scores obtained in Experiment 1 as a reference, we independently replaced one personality dimension, openness, conscientiousness, extraversion, agreeableness, or neuroticism, with an extreme value of either 1 or 5, while keeping the remaining dimensions fixed. The RPD experiments were then conducted under the same settings as in Experiment 2. Figures 3, 4, and 5 present the average cooperation rates, average cumulative payoffs, and their differences for GPT-3.5-turbo, GPT-4o, and GPT-5, respectively. In each figure, panels (A), (B), and (C) correspond to setting the manipulated personality dimension to score = 1, score = 5, and their difference (515-1), respectively.
@@
For GPT-3.5-turbo (Fig. 3), manipulation of agreeableness had the strongest impact on cooperative behavior, followed by extraversion. When agreeableness was set to 1 (A-1) (Fig. 3A), cooperation rates were consistently lower than those observed under other personality manipulations across all opponent strategies, resulting in a large negative difference in average cooperation. In contrast, when agreeableness was set to 5 (A-5) (Fig. 3B), cooperation rates were comparable to those observed under other personality conditions, yielding relatively small differences.
@@
Figure 3: Effects of extreme personality-score manipulations for GPT-3.5-turbo. Left panels show cooperation rates and right panels show cumulative payoffs. (A) Manipulated personality dimension set to score = 1. (B) Manipulated personality dimension set to score = 5. (C) Differences between score = 5 and score = 1 (515-1).
A similar but weaker pattern is observed for extraversion. Under the E-1 condition (Fig. 3A), cooperation rates against ALLD and RANDOM strategies were slightly lower than those under other personality manipulations, which contributed to observable differences in average cooperation. When extraversion was set to 5 (E-5) (Fig. 3B), however, cooperation rates showed little deviation from other conditions.
For GPT-4o (Fig. 4), agreeableness again emerged as the dominant factor influencing cooperative behavior, whereas manipulation of other personality traits had minimal effects. Under the A-1 condition (Fig. 4A), cooperation rates droped to 0% across all opponent strategies, indicating consistently defecting behavior. Under the A-5 condition (Fig. 4B), average cooperation rates increased to levels comparable to or higher than those observed under other personality manipulations. However, cooperation rates against ALLD and RANDOM strategies were highest under A-5, resulting in increased exploitation and, consequently, the lowest average cumulative payoff among all personality conditions.
Figure 4: Effects of extreme personality-score manipulations for GPT-4o. Panel layout and notation are the same as in Fig. 3.
@@
For GPT-5 (Fig. 5), agreeableness similarly had the largest effect on cooperation, while manipulation of other personality traits produces little changed in behavior. Under the A-1 condition (Fig. 5A), cooperation rates fell to 0% across all opponent strategies, mirroring the behavior observed for GPT-4o. Under the A-5 condition (Fig. 5B), cooperation rates increased across strategies; however, the increase in cooperation against ALLD and RANDOM was smaller than that observed for GPT-4o. As a result, GPT-5 experienced less exploitation under high agreeableness than GPT-4o.
Figure 5: Effects of extreme personality-score manipulations for GPT-5. Panel layout and notation are the same as in Fig. 3.
@@ 4 Conclusion
In this study, we investigated how artificially controlled personality traits influenced cooperative behavior in large language model agents. Using the Big Five framework, we first quantitatively measured basic personality scores of multiple LLMs, and then examined their behavior in the RPD games under conditions with and without explicit personality information. We further analyzed the effects of independently manipulating each personality dimension to extreme values.
@@
Overall, our study indicates that personality steering alone is insufficient to fully control cooperative behavior in LLM agents. Rather, cooperative outcomes emerge from the interaction between personality cues and model-specific decision-making processes. This highlights the importance of considering both personality and strategic reasoning when analyzing and designing cooperative AI agents.
Author contributions
M.S., G.I., and W.T. conceptualized the research and designed the methodology. M.S. designed, programmed, and performed the experiments. G.I. supervised the research. M.S. and G.I. wrote the original draft. M.Y. visualized the results. All authors reviewed and edited the manuscript.
Data availability statement
All data are available in the main text or the Supplementary Information.
Appendix. Prompts used in the experiments
This appendix provides the full text of the prompts used in the experiments. All prompts were presented to the agents verbatim, with bracketed variables (e.g., [N], [Action], [X]) dynamically replaced by the actual parameter values used in the experiments.
Appendix A BFI-44 personality questionnaire (as used in [13])
The following prompt corresponds to the Big Five Inventory (BFI-44) used to measure the personality traits of each model. The prompt was presented exactly as shown below, and the agents were instructed to respond by selecting a number from 1 to 5 for each statement. In Experiment 1, this prompt was used to obtain the personality scores of each model.
Here are a number of characteristics that may or may not apply to you.
For example, do you agree that you are someone who likes to spend time
with others? Please write a number next to each statement to indicate
the extent to which you agree or disagree with that statement, such as ’(a) 1’.
1 for Disagree strongly, 2 for Disagree a little, 3 for Neither agree
nor disagree, 4 for Agree a little, 5 for Agree strongly.
(a) Talks a lot
(b) Notices other people’s weak points
(c) Does things carefully and completely
(d) Is sad, depressed
(e) Is original, comes up with new ideas
(f) Keeps their thoughts to themselves
(g) Is helpful and not selfish with others
(h) Can be kind of careless
(i) Is relaxed, handles stress well
(j) Is curious about lots of different things
(k) Has a lot of energy
(l) Starts arguments with others
(m) Is a good, hard worker
(n) Can be tense; not always easy going
(o) Clever; thinks a lot
(p) Makes things exciting
(q) Forgives others easily
(r) Isn’t very organized
(s) Worries a lot
(t) Has a good, active imagination
(u) Tends to be quiet
(v) Usually trusts people
(w) Tends to be lazy
(x) Doesn’t get upset easily; steady
(y) Is creative and inventive
(z) Has a good, strong personality
(aa) Can be cold and distant with others
(ab) Keeps working until things are done
(ac) Can be moody
(ad) Likes artistic and creative experiences
(ae) Is kind of shy
(af) Kind and considerate to almost everyone
(ag) Does things quickly and carefully
(ah) Stays calm in difficult situations
(ai) Likes work that is the same every time
(aj) Is outgoing; likes to be with people
(ak) Is sometimes rude to others
(al) Makes plans and sticks to them
(am) Get nervous easily
(an) Likes to think and play with ideas
(ao) Doesn’t like artistic things (plays, music)
(ap) Likes to cooperate; goes along with others
(aq) Has trouble paying attention
(ar) Knows a lot about art, music and books
[…]
Appendix B Repeated Prisoner’s Dilemma game prompt (based on [26])
This section provides the prompt used to describe the repeated Prisoner’s Dilemma (RPD) game to the agents. The prompt specifies the game context, available actions, payoff structure, and the number of rounds, and was presented verbatim to the agents prior to gameplay.
Context: You and your opponent are playing a multi-round game.
At each turn you and your opponent simultaneously perform one of the
following actions: {"Cooperate", "Defect"}
The payoffs for each combination of chosen actions are the following:
If you play "Cooperate" and your opponent plays "Cooperate", you collect
3 points and your opponent collects 3 points.
If you play "Cooperate" and your opponent plays "Defect", you collect
0 points and your opponent collects 5 points.
If you play "Defect" and your opponent plays "Cooperate", you collect
5 points and your opponent collects 0 points.
If you play "Defect" and your opponent plays "Defect", you collect
1 point and your opponent collects 1 point.
You will play a total of [ITERATIONS] rounds of this game.
Remember that your objective is to get the highest possible amount of
points in the long run.
For all experiments, [ITERATIONS] in this prompt was set to 10.
B.1 History prompt
This subsection describes the prompt used to present the game history to the agents during the RPD game.
B.1.1 First round
This is the first round of the game.
B.1.2 Subsequent rounds
The history of the game in the last [N] rounds is the following:
Round 1: You played "[Action]" and your opponent played "[Action]"
  You collected [X] points and your opponent collected [X] points.
Round 2: You played "[Action]" and your opponent played "[Action]"
  You collected [X] points and your opponent collected [X] points.
In total, you chose "Cooperate" [N] times and chose "Defect" [N] times,
your opponent chose "Cooperate" [N] times and chose "Defect" [N] times.
In total, you collected [X] points and your opponent collected [X] points.
Current round: [N].
In this prompt, [N] refers to either the number of previous rounds or the number of times “Cooperate” or “Defect” was chosen. [Action] corresponds to the action chosen in each round (“Cooperate” or “Defect”), and [X] represents the payoff obtained either in an individual round (0, 1, 3, or 5) or as a cumulative total.
Appendix C Big Five personality trait prompt
This section presents the personality trait prompt provided to each model prior to playing the RPD game in Experiments 2 and 3.
In the baseline condition of Experiment 2, this prompt is not provided, and each model plays the RPD game without any personality-related information. In the personality-informed condition of Experiment 2, the model plays the RPD game after receiving this prompt, which is constructed based on the personality scores obtained in Experiment 1.
In Experiment 3, the personality scores measured in Experiment 1 are also used; however, one personality dimension is independently manipulated to an extreme value, either low (1) or high (5), while the remaining four dimensions are held constant. The model then plays the RPD game under these manipulated personality conditions.
You are an AI agent with the following Big Five personality profile:
[E, A, C, N, O]
Personality Traits:
[Extraversion, Agreeableness, Conscientiousness, Neuroticism, Openness]
Trait Interpretation:
- Extraversion: Lower scores indicate more introverted traits, higher
  scores indicate more extroverted traits
- Agreeableness: Lower scores indicate more antagonistic traits, higher
  scores indicate more agreeable traits
- Conscientiousness: Lower scores indicate more unconscientious traits,
  higher scores indicate more conscientious traits
- Neuroticism: Lower scores indicate more emotionally stable traits,
  higher scores indicate more neurotic traits
- Openness: Lower scores indicate more closed to experience traits,
  higher scores indicate more open to experience traits
Personality scores are on a 1-5 scale. Your traits are described as follows:
- Extraversion (X.X/5.0): [YYY]
- Agreeableness (X.X/5.0): [YYY]
- Conscientiousness (X.X/5.0): [YYY]
- Neuroticism (X.X/5.0): [YYY]
- Openness (X.X/5.0): [YYY]
Decision-Making Guidelines:
- Consider your personality traits when making decisions
- Make choices that feel natural and authentic to your personality profile
- Respond based on how someone with your characteristics might naturally
  approach this situation
Remember: Your personality profile represents your stable characteristics
and tendencies.
In this description, X.X represents the average personality score of each model measured in Experiment 1, while [YYY] refers to the corresponding natural-language explanation based on the score ranges defined in the next subsection.
C.1 Natural-language descriptions of personality score ranges
For each personality dimension, the following natural-language descriptions are used according to the corresponding score range. These descriptions are used to translate numerical personality scores into qualitative statements in the personality trait prompt.
Extraversion:
  • 1.0–1.5: You are highly introverted, strongly preferring solitude and quiet environments over social interactions.
  • 1.5–2.5: You are somewhat introverted, generally preferring solitude but comfortable with limited social interaction.
  • 2.5–3.5: You have a balanced social tendency, comfortable in both social and solitary situations.
  • 3.5–4.5: You are somewhat extraverted, generally seeking social interaction and being energetic in groups.
  • 4.5–5.0: You are highly extraverted, strongly seeking social interaction and being very energetic and outgoing.
Agreeableness:
  • 1.0–1.5: You are highly competitive and skeptical, strongly prioritizing self-interest and being confrontational.
  • 1.5–2.5: You tend to be competitive and skeptical, generally prioritizing self-interest.
  • 2.5–3.5: You balance cooperation and self-advocacy reasonably well.
  • 3.5–4.5: You are generally cooperative and trusting, prioritizing harmony and others’ well-being.
  • 4.5–5.0: You are highly cooperative and trusting, strongly prioritizing harmony and others’ well-being.
Conscientiousness:
  • 1.0–1.5: You are highly spontaneous and flexible, strongly preferring adaptability over rigid planning.
  • 1.5–2.5: You are somewhat spontaneous and flexible, generally preferring adaptability over rigid planning.
  • 2.5–3.5: You balance structure and flexibility in your approach to tasks.
  • 3.5–4.5: You are generally organized and disciplined, preferring structured and systematic approaches.
  • 4.5–5.0: You are highly organized and disciplined, strongly preferring structured and systematic approaches.
Neuroticism:
  • 1.0–1.5: You are highly emotionally stable and resilient, remaining very calm under pressure.
  • 1.5–2.5: You are somewhat emotionally stable and resilient, generally remaining calm under pressure.
  • 2.5–3.5: You have moderate emotional stability with normal stress responses.
  • 3.5–4.5: You are somewhat emotionally sensitive, experiencing worry and stress more frequently.
  • 4.5–5.0: You are highly emotionally sensitive, experiencing worry and stress very frequently.
Openness:
  • 1.0–1.5: You strongly prefer familiar approaches and conventional thinking.
  • 1.5–2.5: You somewhat prefer familiar approaches and conventional thinking.
  • 2.5–3.5: You balance innovation and tradition in your thinking.
  • 3.5–4.5: You are generally open to new experiences, creative, and intellectually curious.
  • 4.5–5.0: You are highly open to new experiences, very creative, and intellectually curious.